OpenStudy (anonymous):

the sum of three consecutive integers is -69 find the three integers

OpenStudy (anonymous):

Let them be x, x+1 and x+2 So x+x+1+x+2=-69 Now solve for x

OpenStudy (wasiqss):

-22 -23 -24

hero (hero):

x + (x+1) + (x+2) = -69 3x + 3 = -69 3x = -72 x = -72/3 x = -24 x+1 = -23 x + 2 = -22

OpenStudy (anonymous):

thanxx

OpenStudy (anonymous):

n + n+1 + n+ 2 = -69 where n is highest number 3n = -72 n = -24 so the3 numbers are -22,-23 -24